# Semantic Segmentation , Self-Driving Car , Computer vision , BDD100K , Light-weight model , Embedded devices

### v4

@@ 4.4 Ablation study
To assess the impact of the multitasking approach on each task, we compare the performance of the TwinLiteNet model in multitasking against its performance in executing each task separately after removing irrelevant decoders. The TwinLiteNet(Drivable only) specializes in segmenting drivable areas, while the TwinLiteNet(Lane only) focuses on lane segmentation. The results, presented in Table 8, show that employing multitask learning not only improves the mIoU for the task of segmenting drivable areas (by an increase of 0.2%), but also leads to certain trade-offs. Specifically, while there’s a slight increase in performance for drivable area segmentation task, the IoU for the lane segmentation task decreases by 0.2% when multitasking, and simultaneously, the model size and computational cost also increase, respectively, by 0.03M and 0.61 GFLOPs.
We then compare the performance of the ESP module and the proposed DESP module for the Encoder block. In the Encoder block (Figure 4), each time the Stride ESP module is executed, a series of DESP modules from the TwinLiteNet+ architecture will be executed instead of the following ESP module, such as ESPNet [13]. To evaluate the effectiveness of replacing the ESP module with the DESP module, we assess the accuracy, computational cost, and model weight and compare the results presented in Table 9. The results indicate negligible differences in mIoU (%) and IoU (%) metrics in the four different configs of TwinLiteNet+. In the Nano and Small configs, there are no changes in FLOPs and parameters when using either the ESP or DESP modules. However, in the Medium and Large configs, both parameters and FLOPs are significantly reduced when using DESP. This highlights the effectiveness of the proposed DESP architecture in reducing computational complexity while maintaining similar performance metrics.
We have undertaken numerous modifications and enhancements, along with extensive experimentation. Table 10 presents a curated list of changes implemented during our experiments and the corresponding improvements integrated into the entire network, culminating in a potent segmentation model for drivable areas and lanes. We construct a Baseline model comprising an Encoder block and a single Decoder block for both tasks, as opposed to two Decoder blocks. The output of this Baseline model is now , encompassing three channels: two for the segmentation tasks and one for the background. Our Baseline does not utilize Partial Class Activation Attention, instead directly concatenating the Encoder and Decoder blocks. During training, this Baseline solely employs Focal Loss for error computation and does not utilize Exponential Moving Average (EMA). Sequentially, we augment the Baseline with Tversky Loss, Multiple Decoder, Partial Class Activation Attention, and EMA. The results demonstrate significant improvements in both segmentation tasks due to our proposed methods. Specifically, with the full implementation of these enhancements, the model exhibited an increase in mIoU for drivable area segmentation from 90.9% to 92.9% and an increase in IoU for lane segmentation from 25.2% to 34.2%. This not only proves the efficacy of each component but also illustrates the power of their combination. Although there is a slight increase in model size and computational cost, these improvements have contributed to setting a new standard in segmentation performance for practical applications.
4.5 Detailed Comparison
In this section, we compare the TwinLiteNet+ model with all models including single-task models and multitask models encompassing segmentation and object detection tasks. In section 4.4, we demonstrate that either single-task or multitask execution could improve accuracy, so it is not surprising that some results indicate our model’s accuracy is lower than some single-task models or those incorporating object detection. Upon examining the numerical data in Table 11, it’s evident that the TwinLiteNet+ models demonstrate a competitive performance across different configurations. While the TwinLiteNet shows modest performance with the lowest parameters, highlighting its potential for lightweight applications, the TwinLiteNet model achieves impressive results, especially in lane segmentation with an IoU of 34.2% and drivable area segmentation with a mIoU of 92.9%, outperforming many other models. This indicates that while our model may not always surpass single-task models in individual tasks, it provides a balanced and efficient solution for multitasking scenarios. Notably, when compared to other models that perform both drivable area and lane segmentation along with vehicle detection, TwinLiteNet shows an impressive balance of high accuracy and low parameter count, suggesting an efficient trade-off between performance and model complexity. These insights underline the efficacy of our model in handling complex, real-world scenarios that require simultaneous perception tasks.
Table 11: More detailed performance comparisons (using mIoU (%) for Drivable Area and IoU (%) metric for Lane Segmentation): The best and second best results are marked in bold and underline respectively.
In this study, we introduce TwinLiteNet+, a novel model that enhances drivable area segmentation and lane detection capabilities and demonstrates efficient deployment on embedded devices. Designed to optimize the balance between accuracy and computational efficiency, the model is configurable from Nano to Large configs to meet diverse requirements. Experimental evaluations reveal that TwinLiteNet+ not only achieves superior accuracy on the BDD100K dataset but also exhibits versatile deployment potential on devices with constrained computational power. While the results are promising, certain challenges still remain, particularly in handling adverse conditions such as harsh weather, low-light or nighttime scenarios. There should be future advanced training techniques to enhance model performance under these challenging conditions, ensuring reliability and robustness in real-world applications. Our proposed model can only perform two main tasks, which is still quite simple; however, it can be further developed to handle additional tasks necessary for self-driving cars, such as object detection and depth estimation. In the end, this work pioneers a novel approach for the development of intelligent driving and driver assistance systems, underscoring the importance of AI model optimization for real-world deployment. By achieving both high accuracy and efficient performance, TwinLiteNet+ paves the way for a new era in AI applications within the autonomous driving domain and contributes to the advancement of safety and efficiency on the roads.

### v5

@@ 4.7 Ablation study
4.7.1 Multi-task versus single-task
To evaluate the effectiveness of the multitask learning strategy, we compare the performance of the TwinLiteNet model in multitask and single-task configurations. In the single-task settings, we retain only the relevant decoder-either for drivable area or lane segmentation-while removing the other to isolate performance. The results are summarized in Table 8.
For the drivable area segmentation task, the multitask version slightly outperforms its single-task counterpart, achieving a 0.3% gain in mIoU. This improvement suggests that shared representations with the lane segmentation task positively influence the learning process by providing complementary structural cues. For the lane segmentation task, multitask learning maintains the same lane accuracy (81.9%) but experiences a slight drop of 0.2% in IoU, which may result from representational interference between the tasks. These findings indicate that while multitask learning introduces minor trade-offs-particularly for tasks sensitive to spatial precision-it also offers measurable advantages in feature sharing and overall model compactness.
In terms of efficiency, the multitask model incurs only a marginal increase in both parameter count (+0.03M) and computational cost (+0.61 GFLOPs) compared to the single-task variants. This slight overhead primarily stems from the addition of a single extra decoder to support the second task, while the encoder is shared between both. In contrast, deploying two separate single-task models would require duplicating the entire backbone, leading to a significantly higher computational cost and memory footprint. Thus, the multitask approach not only enables joint optimization and feature sharing but also offers a more resource-efficient alternative to running independent models. These findings reinforce the practical value of multitask learning for real-time autonomous driving applications, where both accuracy and efficiency are critical.
| Method | Drivable Area | Lane | Parameter | FLOPs | ||||||||||||
| mIoU (%) | Acc (%) | IoU (%) | ||||||||||||||
| Drivable Area (only) | 92.6 | ✘ | ✘ | 1.91M | 16.97G | |||||||||||
| Lane (only) | ✘ | 81.9 | 34.4 | 1.91M | 16.97G | |||||||||||
| Multi-task |
4.7.2 The ablations in different experimental settings
To better understand the contribution of each component in the proposed TwinLiteNet+ architecture, we perform an ablation study based on the TwinLiteNet configuration. Each experiment isolates a specific module to evaluate its impact on segmentation performance and model complexity, with quantitative results summarized in Table 9. Six modified variants are constructed as follows: (1) Tversky Loss: removed by replacing the combined objective with Focal Loss alone for both segmentation branches; (2) DESP: replaced with the original ESP module from the baseline TwinLiteNet; (3) Robust Decoder: disabled by reverting to the simple decoder design of the TwinLiteNet baseline, thus not use the UCB and USB upsampling proposed modules; (4) Multiple Decoder: replaced with a shared single decoder head producing a unified three-channel output for both tasks; (5) PCAA Module: removed by bypassing the attention mechanism and directly linking the encoder output to the decoder; and (6) EMA (Exponential Moving Average): deactivated by using the model’s raw weights instead of the smoothed exponential averages accumulated during training.
Experimental findings show that each component plays an important role in the model’s effectiveness. The removal of Tversky Loss results in a few performance drops in drivable area segmentation. However, it is significant in lane segmentation, highlighting its crucial role in managing class imbalance and improving recall for thin and sparse structures. Substituting DESP with the baseline ESP architecture slightly increases lane segmentation accuracy. However, it introduces a notable rise in FLOPs and parameter count, underscoring the importance of depthwise separable convolution in achieving efficiency. Eliminating any of the Robust Decoder, Multiple Decoder, or PCAA modules leads to consistent degradation across both drivable area and lane segmentation metrics, confirming their importance in enhancing spatial detail, semantic refinement, and task-specific decoding. Lastly, omitting EMA leads to minor reductions in performance, suggesting its stabilizing influence on model convergence and generalization.
The results from the ablation experiments demonstrate that each component in the design of TwinLiteNet+ plays a vital role-not only enhancing accuracy in both segmentation tasks but also optimizing computational cost.
| Drivable Area | Lane | FLOPs | #Param. | |||||||||||||
| mIoU (%) | Acc (%) | IoU (%) | ||||||||||||||
| TwinLiteNet | 92.9 | 81.9 | 34.2 | 17.58G | 1.94M | |||||||||||
| w/o Tversky Loss |
| w/o DESP proposed |
| w/o Robust Decoder |
| w/o Multiple Decoder |
| w/o PCAA Module |
| w/o EMA |
| Model | Drivable Area | Lane | Params |
| mIoU (%) | IoU (%) | ||
| Drivable Area Segmentation (Only) | |||
| PSPNet [7] | 89.6 | ✘ | – |
| MultiNet [8] | 71.6 | ✘ | – |
| R-CNNP [43] | 90.2 | ✘ | – |
| YOLOP [43] | 92.0 | ✘ | – |
| YOLOv8 [61] | 78.1 | ✘ | 03.26M |
| YOLO-D [62] | 92.4 | ✘ | – |
| Lane Segmentation (Only) | |||
| ENet [5] | ✘ | 14.6 | – |
| ENET-SAD [6] | ✘ | 16.0 | – |
| SCNN [9] | ✘ | 15.8 | – |
| YOLOv8segll [61] | ✘ | 22.9 | 03.26M |
| YOLO-L [62] | ✘ | 28.1 | – |
| Drivable Area and Lane Segmentation | |||
| DeepLabV3+ [15] | 90.9 | 29.8 | 015.4M |
| SegFormer [16] | 92.3 | 31.7 | 007.2M |
| R-CNNP [43] | 90.2 | 24.0 | – |
| YOLOP [43] | 91.6 | 26.5 | 05.53M |
| IALaneNetConvNeXt-small [10] | 91.7 | 32.5 | 039.9M |
| YOLOv8 [61] | 84.2 | 24.3 | – |
| Sparse U-PDP [44] | 91.5 | 31.2 | – |
| TwinLiteNet [12] | 91.3 | 31.1 | 00.44M |
| Drivable Area and Lane Segmentation + Vehicle Detection | |||
| HybridNets [20] | 90.5 | 31.6 | 013.8M |
| YOLOP [43] | 91.5 | 26.2 | 007.9M |
| DRMNet [63] | 92.2 | 27.0 | 08.09M |
| YOLOPv2 [64] | 93.2 | 27.3 | 038.9M |
| A-YOLOM(n) [57] | 90.5 | 28.2 | 04.43M |
| A-YOLOM(s) [57] | 91.0 | 28.8 | 13.61M |
| YOLOPX [21] | 93.2 | 27.2 | 032.9M |
| CenterPNets [65] | 92.8 | 32.1 | 028.6M |
| Sparse U-PDP [44] | 92.9 | 32.4 | 12.05M |
| CFFM [66] | 92.8 | 31.1 | - |
| TriLiteNetbase [67] | 92.4 | 29.8 | 02.35M |
| YOLOPv3 [22] | 93.2 | 28.0 | 030.2M |
| TwinLiteNet | 87.3 | 23.3 | 0.03M |
| TwinLiteNet | 90.6 | 29.3 | 0.12M |
| TwinLiteNet | 92.0 | 32.3 | 0.48M |
| TwinLiteNet | 92.9 | 34.2 | 1.94M |
4.8 More comparison
In Table 10, we present a comprehensive comparison of the proposed TwinLiteNet+ model against a diverse set of representative methods selected based on task relevance. These include: (1) single-task models, which focus exclusively on either lane segmentation (e.g., [5, 6, 9, 57, 62]) or drivable area segmentation (e.g., [7, 8, 43, 57, 62]); (2) and models designed explicitly for drivable area and lane segmentation [12, 10, 68, 15, 16], which share the same task formulation as our method; (3) multi-task models that jointly perform drivable area, lane, and object detection, such as [20, 43, 64, 22, 21, 57, 65, 67, 66, 44]. This selection ensures that our evaluation captures related approaches, from task-specific to comprehensive multi-task perception models.
The results indicate that TwinLiteNet+ achieves highly competitive performance, particularly within the lightweight model category. Notably, TwinLiteNet and TwinLiteNet exhibit excellent trade-offs between model size and accuracy, with parameter counts of only 0.03M and 0.12M, respectively-making them well-suited for deployment on resource-constrained embedded systems. While single-task models are typically optimized for a specific objective, they lack the capacity to leverage complementary information across tasks, as discussed in Section 4.7.1. In contrast, the multitask learning strategy employed in TwinLiteNet+ enhances segmentation performance by enabling effective feature sharing, allowing it to outperform dedicated single-task counterparts. On the other hand, multitask models that incorporate object detection often require significantly more complex architectures and higher computational overhead. Thus, a multitasking design focused solely on segmentation tasks-such as TwinLiteNet+-achieves a favorable balance between accuracy, efficiency, and architectural simplicity, making it more suitable for real-world applications on embedded platforms.
While TwinLiteNet achieves a strong mIoU of 92.9% for drivable area segmentation, it falls slightly short of the 93.2% achieved by high-capacity models such as YOLOPv2 [64], YOLOPv3 [22], and YOLOPX [21]. However, these models have a large number of parameters, exceeding 30 million, rendering them less practical for embedded deployment. Nevertheless, TwinLiteNet achieves the highest lane segmentation IoU in the entire benchmark (34.2%), outperforming the second-best IALaneNet [10] (32.5%), despite IALaneNet requiring over 39.9M parameters-more than 20 times that of TwinLiteNet.
These results provide a broader perspective by evaluating TwinLiteNet+ against both single-task and multi-task approaches, including those that include object detection, further ensuring the effectiveness and scalability of our proposed lightweight model.
5 Discussion and Conclusion
5.1 Limitations and future work
Although TwinLiteNet+ achieves a reasonable balance between accuracy and computational efficiency, there are still several limitations that should be addressed in future research. First, the model is currently trained and evaluated under general environmental conditions and has not been optimized for specific scenarios. As a result, its performance tends to degrade in adverse weather conditions or cases with limited training samples. Second, TwinLiteNet+ currently focuses on only two perception tasks: drivable area segmentation and lane detection. While this design helps maintain a compact and efficient architecture, it lacks the comprehensiveness required by more advanced autonomous driving systems, which typically demand simultaneous handling of multiple tasks such as object detection, depth estimation, or traffic sign recognition. Third, although the model is lightweight and efficient, it relies on manually designed architectural configurations. The lack of neural architecture search (NAS) techniques limits its adaptability to diverse deployment requirements in real-world settings.
Addressing these limitations will be essential further to enhance the model’s performance and practical applicability. Doing so will not only improve TwinLiteNet+ itself but also contribute to the advancement of more intelligent, efficient, and reliable multi-task perception systems for autonomous driving.
5.2 Conclusion
In this study, we presented TwinLiteNet+, a novel and efficient multi-task segmentation model designed for drivable area and lane detection. By enhancing both the encoder and decoder architecture while maintaining a lightweight design, TwinLiteNet+ achieves a strong balance between segmentation accuracy and computational efficiency. The model is offered in four scalable configurations, making it adaptable to a wide range of hardware-from high-performance GPUs to resource-constrained embedded systems. Extensive experiments on the BDD100K dataset demonstrate that TwinLiteNet+ not only delivers competitive accuracy compared to existing methods but also ensures real-time performance suitable for practical deployment. These results highlight the model’s potential as a reliable perception module in intelligent driving systems, where both efficiency and accuracy are critical. This work contributes a promising solution for real-world autonomous driving applications, where streamlined architectures and performance-aware designs are increasingly essential.
6 Acknowledgement
This research was supported by The VNUHCM-University of Information Technology’s Scientific Research Support Fund.